OpenStudy (anonymous):

Allen has four boxes. Each box contains four blue balls and two yellow balls. What is the probability that he will choose a yellow ball from each box?

OpenStudy (anonymous):

4 out of 8

OpenStudy (anonymous):

am I right

OpenStudy (anonymous):

okay but I mean WHYYYY Also the answer key says 1 out of 81 but we're not quite sure why...?

OpenStudy (anonymous):

8 out of 16

OpenStudy (anonymous):

oh sorry wait answer key then why do you need help

OpenStudy (anonymous):

\[\left(\frac{2}{6}\right)^4=\frac{1}{81} \]

OpenStudy (anonymous):

Because we aren't sure how to get that answer

OpenStudy (anonymous):

THANK YOU ROBTOBEY

OpenStudy (anonymous):

Your welcome.